The Eiffel Tower is 497 ft. taller than he Washington Monument. If each of the monuments were 58ft shorter, the Eiffel tower would be twice as tall as the Washington Monument. How tall is each?
Answer by mananth(16949)

WM be x
ET = x+497
..
if each were shorter by 58 ft
their height would be
WM x-58
ET x+497-58
...
x+497-58 = 2(x-58)
x+439=2x-116
2x-x=439+116
x= 555 feet the washington Monumnet (WM)
ET = x+497
=555+497=1052 feet
...
m.ananth@hotmail.ca